# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
$F_{R}(\nu)\coloneqq(R_{\mathrm{tr}}+\nu I_{d})^{-1},$ (SI.8.1)
where $R_{\mathrm{tr}}$ is the sample covariance matrix of the task vectors as defined in (SI.4.1) and $\nu$ is a positive scalar.
@@
$\frac{1}{d}\operatorname{tr}F_{R}(\nu)\simeq\mathcal{M}_{\kappa}(\nu)$ (SI.8.2)
@@
$\mathcal{M}_{\kappa}(\nu)\coloneqq\frac{2}{\nu+1-1/\kappa+\left[(\nu+1-1/\kappa)^{2}+4\nu/\kappa\right]^{1/2}}.$ (SI.8.3)
@@
$\frac{1}{\mathcal{M}_{\kappa}(\nu)}=\frac{1}{1+\mathcal{M}_{\kappa}(\nu)/\kappa}+\nu.$ (SI.8.4)
@@
$\frac{1}{d}\operatorname{tr}F^{2}(\nu)\simeq-\mathcal{M}^{\prime}_{\kappa}(\nu)=\frac{\mathcal{M}^{2}_{\kappa}(\nu)}{1-\frac{\kappa\mathcal{M}_{\kappa}^{2}(\nu)}{[\kappa+\mathcal{M}_{\kappa}(\nu)]^{2}}}.$ (SI.8.5)
For the remainder of this appendix, we will further explore the self-consistent equation given by (SI.4.7). We will show that the solution $\chi_{\pi}$ and its derivative $\frac{d}{d\pi}\chi_{\pi}$, at $\pi=0$, can be characterized by the function $\mathcal{M}_{\kappa}(\nu)$ in (SI.8.3). To start, note that at $\pi=0$, the equation in (SI.4.7) can be written as
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}.$ (SI.8.6)
Recall the definition of $E_{\mathrm{tr}}$ given in (SI.4.3). It is straightforward to verify that
$(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I_{d+1})^{-1}=\begin{bmatrix}F_{R}(\nu_{0})+a^{\ast}(1+\rho)^{2}F_{R}(\nu_{0})b_{\mathrm{tr}}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&-a^{\ast}(1+\rho)F_{R}(\nu_{0})b_{\mathrm{tr}}\\ -a^{\ast}(1+\rho)b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&a^{\ast}\end{bmatrix},$ (SI.8.7)
where $F_{R}(\cdot)$ is the function defined in (SI.8.1),
$\nu_{0}=\frac{1+\rho}{\alpha}+\lambda(1+\chi_{0})$ (SI.8.8)
@@
$\frac{1}{a^{\ast}}=(1+\rho)^{2}+\lambda(1+\chi_{0})-(1+\rho)^{2}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}.$ (SI.8.9)
From (SI.8.7), the equation (SI.8.6) becomes
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}F_{R}(\nu_{0})-(1+\rho)^{2}\frac{a^{\ast}\lambda(1+\chi_{0})}{d}\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\rVert^{2}.$ (SI.8.10)
@@
$b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}\leq 1\qquad\text{and}\qquad\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\rVert^{2}\leq\frac{1}{\nu_{0}}\leq\frac{\alpha}{1+\rho}.$ (SI.8.11)
Substituting the first inequality above into (SI.8.9) gives us
$a^{\ast}\lambda(1+\chi_{0})\leq 1.$ (SI.8.12)
Combining this estimate with the second inequality in (SI.8.11), we can conclude that the last term on the right-hand side of (SI.8.10) is negligible as $d\to\infty$. Moreover, using the asymptotic characterization given in (SI.8.2), the equation (SI.8.10) leads to
$\frac{\tau\chi_{0}}{1+\chi_{0}}\simeq 1-\lambda(1+\chi_{0})\mathcal{M}_{\kappa}(\nu_{0}).$ (SI.8.13)
@@
$\xi_{0}=\lambda(1+\chi_{0}),$ (SI.8.14)
and also recalling the definition of $\nu_{0}$ in (SI.8.8), we can further transform (SI.8.13) to
$\xi_{0}\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi_{0}\right)-\frac{\tau\lambda}{\xi_{0}}\simeq 1-\tau.$ (SI.8.15)
Observe that the above is identical to the equation in (SI.5.6), except for a small error term captured by $\simeq$. By the stability of (SI.5.6), we can then conclude that
$\xi_{0}\simeq\xi,$ (SI.8.16)
thus verifying (SI.5.16).
Next, we compute $\chi^{\prime}_{0}$, the derivative of $\chi_{\pi}$ (with respect to $\pi$) evaluated at $\pi=0$. Differentiating (SI.4.7) give us
$\tau\chi^{\prime}_{0}=\frac{1}{d}\operatorname{tr}\left[(E_{\mathrm{tr}}+\xi_{0}I)^{-1}\Big{(}\chi^{\prime}_{0}E_{\mathrm{tr}}-\frac{(1+\chi_{0})^{2}}{\tau}B_{\mathrm{test}}\Big{)}(E_{\mathrm{tr}}+\xi_{0}I)^{-1}E_{\mathrm{tr}}\right].$ (SI.8.17)
@@
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}/d+\xi^{2}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}/d-\tau},$ (SI.8.18)
where we have used (SI.8.16) to replace $\xi_{0}$ in (SI.8.17) by $\xi$, with the latter being the solution to the self-consistent equation in (SI.5.6). Using the decomposition in (SI.8.7) and following similar arguments that allowed us to simplify (SI.8.10) to (SI.8.13), we can check that
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (SI.8.19)
@@
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-2}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F^{2}\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq-\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (SI.8.20)
where $S$ is a $d\times(d+1)$ matrix obtained by removing the last row of $I_{d+1}$, and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (SI.8.3). Substituting (SI.8.19) and (SI.8.20) into (SI.8.18) yields
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\tau}.$ (SI.8.21)

### v3

@@ Appendix H The Stieltjes Transforms of Wishart Ensembles
$F_{R}(\nu)\coloneqq(R_{\mathrm{tr}}+\nu I_{d})^{-1},$ (250)
where $R_{\mathrm{tr}}$ is the sample covariance matrix of the task vectors as defined in (100) and $\nu$ is a positive scalar.
@@
$\frac{1}{d}\operatorname{tr}F_{R}(\nu)\simeq\mathcal{M}_{\kappa}(\nu)$ (251)
@@
$\mathcal{M}_{\kappa}(\nu)\coloneqq\frac{2}{\nu+1-1/\kappa+\left[(\nu+1-1/\kappa)^{2}+4\nu/\kappa\right]^{1/2}}.$ (252)
@@
$\frac{1}{\mathcal{M}_{\kappa}(\nu)}=\frac{1}{1+\mathcal{M}_{\kappa}(\nu)/\kappa}+\nu.$ (253)
@@
$\frac{1}{d}\operatorname{tr}F^{2}(\nu)\simeq-\mathcal{M}^{\prime}_{\kappa}(\nu)=\frac{\mathcal{M}^{2}_{\kappa}(\nu)}{1-\frac{\kappa\mathcal{M}_{\kappa}^{2}(\nu)}{[\kappa+\mathcal{M}_{\kappa}(\nu)]^{2}}}.$ (254)
For the remainder of this appendix, we will further explore the self-consistent equation given by (106). We will show that the solution $\chi_{\pi}$ and its derivative $\frac{d}{d\pi}\chi_{\pi}$, at $\pi=0$, can be characterized by the function $\mathcal{M}_{\kappa}(\nu)$ in (252). To start, note that at $\pi=0$, the equation in (106) can be written as
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}.$ (255)
Recall the definition of $E_{\mathrm{tr}}$ given in (102). It is straightforward to verify that
$(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I_{d+1})^{-1}=\begin{bmatrix}F_{R}(\nu_{0})+a^{\ast}(1+\rho)^{2}F_{R}(\nu_{0})b_{\mathrm{tr}}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&-a^{\ast}(1+\rho)F_{R}(\nu_{0})b_{\mathrm{tr}}\\ -a^{\ast}(1+\rho)b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&a^{\ast}\end{bmatrix}$, (256)
where $F_{R}(\cdot)$ is the function defined in (250),
$\nu_{0}=\frac{1+\rho}{\alpha}+\lambda(1+\chi_{0})$ (257)
@@
$\frac{1}{a^{\ast}}=(1+\rho)^{2}+\lambda(1+\chi_{0})-(1+\rho)^{2}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}.$ (258)
From (256), the equation (255) becomes
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}F_{R}(\nu_{0})-(1+\rho)^{2}\frac{a^{\ast}\lambda(1+\chi_{0})}{d}\mathinner{\!\left\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\right\rVert}^{2}.$ (259)
@@
$b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}\leq 1\qquad\text{and}\qquad\mathinner{\!\left\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\right\rVert}^{2}\leq\frac{1}{\nu_{0}}\leq\frac{\alpha}{1+\rho}.$ (260)
Substituting the first inequality above into (258) gives us
$a^{\ast}\lambda(1+\chi_{0})\leq 1.$ (261)
Combining this estimate with the second inequality in (260), we can conclude that the last term on the right-hand side of (259) is negligible as $d\to\infty$. Moreover, using the asymptotic characterization given in (251), the equation (259) leads to
$\frac{\tau\chi_{0}}{1+\chi_{0}}\simeq 1-\lambda(1+\chi_{0})\mathcal{M}_{\kappa}(\nu_{0}).$ (262)
@@
$\xi_{0}=\lambda(1+\chi_{0}),$ (263)
and also recalling the definition of $\nu_{0}$ in (257), we can further transform (262) to
$\xi_{0}\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi_{0}\right)-\frac{\tau\lambda}{\xi_{0}}\simeq 1-\tau.$ (264)
Observe that the above is identical to the equation in (152), except for a small error term captured by $\simeq$. By the stability of (152), we can then conclude that
$\xi_{0}\simeq\xi,$ (265)
thus verifying (162).
Next, we compute $\chi^{\prime}_{0}$, the derivative of $\chi_{\pi}$ (with respect to $\pi$) evaluated at $\pi=0$. Differentiating (106) give us
$\tau\chi^{\prime}_{0}=\frac{1}{d}\operatorname{tr}\left[(E_{\mathrm{tr}}+\xi_{0}I)^{-1}\Big(\chi^{\prime}_{0}E_{\mathrm{tr}}-\frac{(1+\chi_{0})^{2}}{\tau}B_{\mathrm{test}}\Big)(E_{\mathrm{tr}}+\xi_{0}I)^{-1}E_{\mathrm{tr}}\right].$ (266)
@@
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}/d+\xi^{2}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}/d-\tau},$ (267)
where we have used (265) to replace $\xi_{0}$ in (266) by $\xi$, with the latter being the solution to the self-consistent equation in (152). Using the decomposition in (256) and following similar arguments that allowed us to simplify (259) to (262), we can check that
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-1}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (268)
@@
$\frac{1}{d}\operatorname{tr}(E_{\mathrm{tr}}+\xi I)^{-2}\simeq\frac{1}{d}\operatorname{tr}S(E_{\mathrm{tr}}+\xi I)^{-2}S^{\top}\simeq\frac{1}{d}\operatorname{tr}F^{2}\left(\frac{1+\rho}{\alpha}+\xi\right)\simeq-\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right),$ (269)
where $S$ is a $d\times(d+1)$ matrix obtained by removing the last row of $I_{d+1}$, and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (252). Substituting (268) and (269) into (267) yields
$\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}$ $\simeq\frac{\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}]\right)}{1-2\xi\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\tau}.$ (270)